### v1

@@ 4.3 Tensile Bar
In Fig. 7 we visually compare the extracted Jacobi sets in the original data as seen in the upper Fig. 6(a)–h with the datasets after applying the CA variant A as seen in the Fig. 6(i)–p. The range area is displayed using color coding, with the saturation indicating how high the stresses in the dataset are. The first thing to recognize is that the Jacobi sets in all datasets are visually simplified after applying the CA variant A and parts with small Jacobi set components could be greatly reduced. Differences can be seen in the datasets with notches and the datasets with holes. For example, Fig. 6(a), i shows that the Jacobi sets are simplified at the notches and the structure becomes more recognizable for domain experts. In Fig. 6(c), k, it can be seen that in parts where a high density of small Jacobi set components occurs, the Jacobi sets are simplified, but the symmetry of the dataset is partially broken. For example, in Fig. 6(e), m, it can be seen that the Jacobi sets are simplified in the part around the hole. In Fig. 6(f), n it can also be seen that the Jacobi sets can be simplified at the upper and lower end of the tie rod where low stress values also appear in the dataset, which can also be seen from the white color. In the Section 4 it can be seen that CA variant A can greatly reduce all two measures. The results of the loop subdivision can also be seen here. It is noticeable that the length of all Jacobi sets increases significantly compared to the original data. For the smoothing filters, we have no results for these datasets, as the used filters only work on structured grids. Overall, you can see in the figures that the noise in the datasets can be greatly reduced. The simplification from the Jacobi set components for the most complex tensile bar D from $519$ to $40$ using the CA variant A takes $0.085$ s. All the other tensile bars need less time.
@@ 4.4 Hurricane Isabel
Another real-world dataset from the field of meteorology is Hurricane Isabel[41, 27], which was published as a freely available dataset as part of the 2004 SciVis contest. This modeled hurricane is based on Hurricane Isabel which caused severe destruction in September 2003. The dataset is a 3D dataset with a resolution of $500\times 500\times 100$ and $48$ time steps and contains $13$ scalar variables, such as precipitation, at each data point. For our comparison, we use a layer from the dataset at height $50$ at time step $30$ and triangulate the grid.
@@
To do this, we look at the scalar fields for pressure and temperature for the corresponding visual analysis in Fig. 8. This figure shows the algorithms to be compared, and the original data. An overall view of the dataset can be seen in Fig. 7(c), 7(d), 7(i), and Fig. 7(j). Visually, it can be seen that all $3$ algorithms simplify the Jacobi sets compared to the original data. In Fig. 7(a), 7(e), 7(g) and Fig. 7(k) a cutout of the center of the hurricane is shown. It is easy to see that CA variant A and the loop subdivision simplify the Jacobi sets but still retain the structures. In contrast, the Gaussian filter simplifies the center so much that it looks completely different compared to the original data. A visual comparison in Fig. 7(j) shows that the Gaussian filter greatly simplifies many large structures. A second cutout of the edge of the hurricane can be seen in Fig. 7(b), 7(f), 7(h), and Fig. 7(l). Here it can be seen in the original data that in the parts where the difference between the range areas is small, many small Jacobi set components occur here. Visually, these small Jacobi set components have decreased in the loop subdivision as well as in the CA variant A. This also shows that the Gaussian filter greases the data too much. The Section 4 clearly shows that CA variant A reduces the Jacobi set components the most. It can also be seen that the loop subdivision leads to more Jacobi set components despite a visual simplification. For the Gaussian filter, it can be seen that it reduces the components of the Jacobian sets and even reduces the length of the Jacobian sets the most, but this is at the expense of important structures.
Simplifying the Jacobi set components from $34443^{\prime}$ 344 to $6572^{\prime}$ 657 using the CA variant A takes $5.5$ s.
@@ 5 Discussion
In this study, we compared the variants of the collapse algorithm, smoothing filters, and a loop subdivision with the original data using different data sets to investigate the effect on the Jacobi set simplification. In addition, we visually examined the influence of the different neighborhood graphs on the result using two measures and compared the best variant with the smoothing filters and the loop subdivision.
@@
Our investigations have shown that the structure of the neighborhood graph is of great importance, as the results differed from one another. The idea of converting the neighborhood graph into a neighborhood tree had certain advantages, as this made it possible to assign an orientation direction in particularly noisy parts. However, the decision as to whether the negative (variant B) or the positive orientation (variant C) was preferred led to very different results. Therefore, the choice of variant varies depending on the case. The assignment of the cells according to their local neighborhood (variant D) could already solve this problem better but led to the fact that many small Jacobi set components were still not simplified because they were assigned according to their neighborhood. For this reason, the neighborhood graph was used in collapse algorithm variant A without further adjustments. This has the advantage that the structure is simpler and in the examples considered, the results were not worse, but even better. However, there were small noisy parts in Topological Simplifcation of Jacobi Sets for Piecewise-Linear Bivariate 2D Scalar Fields with Adjustment of the Underlying Data at the bottom left where not all the noise could be removed. Compared to the original data, the results of the collapse algorithm variant A are very good. The neighborhood graphs can be greatly simplified, which is particularly beneficial for domain experts, as the analysis becomes more difficult as the data gets larger and larger. This can also be seen when looking at all datasets together, whereby collapse algorithm variant A was able to reduce the number of Jacobi set components by more than an order of magnitude on average compared to the original data. The length of the Jacobi sets and the number of Jacobi set components were reduced by more than half on average with collapse algorithm variant A. The visual results also show that the dataset provides a significantly better overview. The comparison of the collapse algorithm variant A with the loop subdivision showed that the collapse algorithm does not affect the essential structures in the dataset and can simplify the datasets at the same time. However, there were still deficits, particularly in the case of data with high symmetry and very noisy parts. Here, the collapse algorithm should be further optimized to deliver even better results. Although the loop subdivision can visually provide very good results, the actual problem of noisy small Jacobi set components is not solved, but the triangles where this occurs become smaller, as already shown in the work of Klötzel et al. [27] can be seen. The comparison with the smoothing filters, in particular the Gaussian filter, showed that although it removes common noise in the green regions of interest very well, the essential structures are also changed and thus the content is lost. This can also be seen in Fig. 7(k) where the Gaussian filter can simplify very noisy parts, leaving many small Jacobi set components. Further comparative tests with different $\sigma$’s for the Gaussian filter can be found in the appendix.
@@
Overall, it can be seen that the collapse algorithm variant A can greatly simplify the Jacobi sets without removing essential structures and achieves better results compared to the other algorithms. Also, the time analysis shows that this depends on the number of Jacobi set components to be reduced. On average, the CA variant A requires $13.80$ ms for all datasets to reduce $100$ Jacobi set components.
@@ 6 Conclusion & Future Work
In this study, an algorithm for simplifying Jacobi sets for bivariate 2D scalar fields in unstructured triangulated grids was presented by adjusting the underlying data. This algorithm is based on the collapsing of cells which changes the underlying data, whereby the collapsing cells are identified using a neighborhood graph. This resulted in four algorithm variants. How well the algorithm variants, smoothing filters, and loop subdivision can simplify the Jacobi sets was then studied with the original data using three datasets from the literature. The best of the four variants was determined, which is the collapse algorithm variant A, especially with regard to the number of components of the Jacobi set, which was an order of magnitude lower than in the original data. The Gaussian filter, on the other hand, was able to reduce the length of the Jacobi sets the most, but changed the data set too much. However, the collapse algorithm was also able to reduce the length of the Jacobi sets by half. The results of the collapse algorithm are promising, although some adjustments are still needed for symmetric data such as the tension rods to better preserve symmetry. In addition, an adapted Jacobi set visualization was presented to account for degenerate cells and assign them to a Jacobi set component according to their neighborhood.

### v2

@@ 4.3 Tensile Bar
In Fig. 7 we visually compare the extracted Jacobi sets in the original data as seen in the upper Fig. 6(a)–h with the datasets after applying the CA variant A as seen in the Fig. 6(i)–p. The range area is displayed using color coding, with the saturation indicating how high the stresses in the dataset are. The first thing to recognize is that the Jacobi sets in all datasets are visually simplified after applying the CA variant A and parts with small Jacobi set components could be greatly reduced. Differences can be seen in the datasets with notches and the datasets with holes. For example, Fig. 6(a), i shows that the Jacobi sets are simplified at the notches and the structure becomes more recognizable for domain experts. In Fig. 6(c), k, it can be seen that in parts where a high density of small Jacobi set components occurs, the Jacobi sets are simplified, but the symmetry of the dataset is partially broken. For example, in Fig. 6(e), m, it can be seen that the Jacobi sets are simplified in the part around the hole. In Fig. 6(f), n it can also be seen that the Jacobi sets can be simplified at the upper and lower end of the tie rod where low stress values also appear in the dataset, which can also be seen from the white color. In the Section 4 it can be seen that CA variant A can greatly reduce all two measures. The results of the loop subdivision can also be seen here. It is noticeable that the length of all Jacobi sets increases significantly compared to the original data. For the smoothing filters, we have no results for these datasets, as the used filters only work on structured grids. Overall, you can see in the figures that the noise in the datasets can be greatly reduced. The simplification from the Jacobi set components for the tensile bar D from $519$ to $40$ using the CA variant A takes $0.085$ s. All the other tensile bars need less time.
@@ 4.4 Hurricane Isabel
Another real-world dataset from the field of meteorology is Hurricane Isabel[41, 27] , which was published as a freely available dataset as part of the 2004 SciVis contest. This modeled hurricane is based on Hurricane Isabel which caused severe destruction in September 2003. It is a 3D dataset with a resolution of $500\times 500\times 100$ and $48$ time steps and contains $13$ scalar variables, such as velocity, at each data point. We use a layer from the dataset at height $50$ at time step $30$ and triangulate the grid.
@@
To do this, we look at the scalar fields for pressure and temperature for the corresponding visual analysis in Fig. 8. This figure shows the algorithms to be compared, and the original data. An overall view of the dataset can be seen in Fig. 7(c), 7(d), 7(i), and Fig. 7(j). Visually, it can be seen that all $3$ algorithms simplify the Jacobi sets compared to the original data. In Fig. 7(a), 7(e), 7(g) and Fig. 7(k) a cutout of the center of the hurricane is shown. It is easy to see that CA variant A and the loop subdivision simplify the Jacobi sets but still retain the structures. In contrast, the Gaussian filter simplifies the center so much that it looks completely different compared to the original data. A visual comparison in Fig. 7(j) shows that the Gaussian filter greatly simplifies many large structures. A second cutout of the edge of the hurricane can be seen in Fig. 7(b), 7(f), 7(h), and Fig. 7(l). Here it can be seen in the original data that many small Jacobi set components occur in the part where the difference between the range areas is small. Visually, these small Jacobi set components have decreased in the loop subdivision as well as in the CA variant A. This also shows that the Gaussian filter greases the data too much. Section 4 clearly shows that the CA variant A reduces the components of the Jacobi set the most and that the loop subdivision leads to more Jacobi set components despite an optical simplification. For the Gaussian filter, it can be seen that it reduces the components of the Jacobi sets and even reduces the length of the Jacobi sets the most, but this is at the expense of important structures. Simplifying the Jacobi set components from $34443^{\prime}$ 344 to $6572^{\prime}$ 657 using the CA variant A takes $5.5$ s.
@@ 5 Discussion
In this study, we compared the collapse algorithm (CA) variants, smoothing filters, and a loop subdivision with the original data using different datasets to investigate the effect on the Jacobi set simplification. Besides, we visually examined the influence of the different neighborhood graphs on the result and compared the best variant with the smoothing filters and the loop subdivision.
@@
Our investigations have shown that the structure of the neighborhood graph is of great importance, as the results differed from one another. The idea of converting the neighborhood graph into a neighborhood tree had certain advantages, as this made it possible to assign an orientation direction in particularly noisy parts. However, the decision as to whether the negative (variant B) or the positive orientation (variant C) was preferred led to very different results. Therefore, the choice of variant varies depending on the case. The assignment of the cells according to their local neighborhood (variant D) could already solve this problem better but led to the fact that many small Jacobi set components were still not simplified because they were assigned according to their neighborhood. For this reason, the neighborhood graph was used in CA variant A without further adjustments. This has the advantage that the structure is simpler and in the examples considered, the results were not worse, but even better. However, there were small noisy parts in Topological Simplifcation of Jacobi Sets for Piecewise-Linear Bivariate 2D Scalar Fields with Adjustment of the Underlying Data at the bottom left where not all the noise could be removed. Compared to the original data, the results of the CA variant A are very good. The neighborhood graphs can be greatly simplified, which is particularly beneficial for domain experts, as the analysis becomes more difficult as the data gets larger and larger. This can also be seen when looking at all datasets together, whereby CA variant A was able to reduce the number of Jacobi set components by more than an order of magnitude on average compared to the original data. The length of the Jacobi sets and the number of Jacobi set components were reduced by more than half on average with CA variant A. The visual results also show that the dataset provides a significantly better overview. The comparison of the CA variant A with the loop subdivision showed that the CA does not affect the essential structures in the dataset and can simplify the datasets at the same time. However, there were still deficits, particularly in the case of data with high symmetry and very noisy parts. Here, the CA should be further optimized to deliver even better results. Although the loop subdivision can visually provide very good results, the actual problem of noisy small Jacobi set components is not solved, but the triangles where this occurs become smaller, as already shown in the work of Klötzel et al. [27] can be seen. The comparison with the Gaussian filter, showed that although it removes common noise in the green regions of interest very well, the essential structures are also changed and thus the content is lost. This can also be seen in Fig. 7(k) where the Gaussian filter can simplify very noisy parts, leaving many small Jacobi set components. Further comparative tests with different smoothing filters can be found in the appendix.
@@
Overall, it can be seen that the CA variant A can greatly simplify the Jacobi sets without removing essential structures and achieves better results compared to the other algorithms. Since the algorithm collapses cells, this can lead to the fact that a component cannot be completely removed if the algorithm runs into an oscillation, and at this point, the neighborhood graph cannot be simplified further. Examining the runtime in all datasets, it can be seen that for CA variant A it depends on the number of Jacobi set components to be reduced and requires an average of $13.80$ṁs to reduce $100$ Jacobi set components.
@@ 6 Conclusion & Future Work
In this study, an algorithm for simplifying Jacobi sets for bivariate 2D scalar fields in unstructured triangulated grids was presented by adjusting the underlying data. This algorithm is based on the collapsing of cells which changes the underlying data, whereby the collapsing cells are identified using a neighborhood graph. This resulted in four algorithm variants. How well the algorithm variants, smoothing filters, and loop subdivision can simplify the Jacobi sets was then studied with the original data using three datasets from the literature. The best of the four variants was determined, which is the collapse algorithm variant A, especially with regard to the number of components of the Jacobi set, which was an order of magnitude lower than in the original data. The Gaussian filter, on the other hand, was able to reduce the length of the Jacobi sets the most, but changed the dataset too much. However, the collapse algorithm was also able to reduce the length of the Jacobi sets by half. The results of the collapse algorithm are promising, although some adjustments are still needed for symmetric data such as the tension rods to better preserve symmetry. Besides, an adapted Jacobi set visualization was presented to account for degenerate cells and assign them to a Jacobi set component according to their neighborhood.